# Co-AttenDWG: Co-Attentive Dimension-Wise Gating and Expert Fusion for Multi-Modal Offensive Content Detection

### v1

@@ III-E Classification
TABLE III: Ablation Study and Full Model Performance on Memotion and MIMIC Datasets (Accuracy and Macro F1 in %)
Experiment MIMIC (%) Memotion (%)
Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$)
No ExpertFusion 83.58 83.56 80.90 80.84
No Co-Attention 85.59 85.57 81.19 81.01
No Cross-Attention 83.58 83.57 79.12 79.26
No MambaFormer 84.98 84.95 81.19 81.31
No Conv in MambaFormer 84.78 84.77 79.49 79.63
Co-AttenDWG 87.19 87.16 84.29 84.26
@@
TABLE IV: Co-AttenDWG Performance using different backbone models on MIMIC and Memotion Datasets (Accuracy and F1 in %)
Backbone Datasets
MIMIC (%) Memotion (%)
Text Image Acc F1 Acc F1
mBERT ResNet50 85.29 85.26
mBERT VGG16 86.77 86.83
mBERT EfficientNetV2 85.28 85.27
XLM-RoBERTa ResNet50 87.79 87.83
XLM-RoBERTa VGG16 86.17 86.82
XLM-RoBERTa EfficientNetV2 82.58 82.86
BERT ResNet50 84.29 84.26
BERT VGG16 83.66 83.21
BERT EfficientNetV2 83.66 83.21
DistilBERT ResNet50 81.81 81.89
DistilBERT VGG16 82.09 82.97
DistilBERT EfficientNetV2 79.97 80.01
@@
(a) Confusion matrix for the MIMIC dataset. The true classes are Non Misogynistic (top-left) and Misogynistic (bottom-right).
(b) Confusion matrix for the SemEval Memotion 1.0 dataset, showing not_offensive (top-left), slight, very_offensive, and hateful_offensive classes (bottom-right).
Figure 6: Confusion matrices for (a) the MIMIC dataset and (b) the Memotion dataset. The diagonal cells represent correct classifications, while off-diagonal cells indicate misclassifications. The color intensity in each cell corresponds to the proportion of samples relative to the total in that true class.
IV Experiment and Result Analysis
IV-A Datasets
We evaluate our approach using two publicly available datasets. The first is the SemEval-2020 Memotion Analysis 1.0 dataset [38], which serves as a benchmark for offensive content detection in social media and provides diverse examples of memes along with detailed annotations that capture varying degrees of offensiveness. The second dataset, MIMIC: Misogyny Identification in Multimodal Internet Content in Hindi-English Code-Mixed Language [39], focuses specifically on detecting misogynistic content within multimodal posts characterized by code-mixed language, posing unique challenges for cross-modal understanding. Together, these datasets offer a comprehensive testing ground for our Co-AttenDWG architecture, enabling us to assess its performance in diverse and culturally nuanced contexts. Table I presents dataset statistics.
IV-B Implementation Details
We implement our model in Python 3.12.1 using PyTorch version 2.0.1 on an NVIDIA RTX 2060 GPU with 16 GB of RAM. We use the AdamW optimizer with a fixed learning rate of $2\times 10^{-5}$ and train the model for 20 epochs while employing a learning rate scheduler that dynamically adjusts the rate during training. In our architecture, we set the number of attention heads in the fusion modules to 8 and use 4 heads in the additional self-attention refinement layer. The MambaFormer encoders are configured with a kernel size of 3, a depth of 2 layers, and a dropout rate of 0.1 across all modules. We fine-tune all components, including the pre-trained text encoder (BERT, XLM-RoBERTa) and the image encoder (ResNet50), in an end-to-end manner.
For data preparation, we clean and normalize the input text and images, tokenize the text using the BERT tokenizer, and resize and normalize the images to the required input dimensions. Specifically, for the MIMIC dataset, images are resized to 200 $\times$ 200 pixels. For the SemEval Memotion 1.0 dataset, we use a lower resolution of 160 $\times$ 160 pixels due to resource constraints. We split our datasets from MIMIC and SemEval Memotion 1.0 into 80% for training and 20% for testing and evaluate our model using test accuracy and macro F1-score as our primary metrics. To further enhance performance, we address class imbalance through upsampling [40], ensuring that minority classes are well represented during training. These specific hyperparameter settings, along with our rigorous data preparation and evaluation protocols, enable our model to effectively learn cross-modal interactions and robustly detect offensive content in multimodal data.
@@ IV-C Baseline Comparison
Table II presents a comprehensive comparison of baseline model performance on the MIMIC and SemEval Memotion 1.0 datasets. The table reports two key metrics, Test Accuracy and Macro F1 (both expressed as percentages, with higher values indicating better performance), for each model. We compare several unimodal models (for example, mBERT, BERT, and DistilBERT for text and VGG16, ResNet50, EfficientNetV2, and XLM-RoBERTa for image) as well as various multimodal models that combine different text and image backbones in diverse configurations. Notably, our proposed model, Co-AttenDWG, is listed at the bottom of the table to highlight its superior performance. For clarity, the second-best scores in each column are underlined (for instance, VisualBERT and mCLIP achieve the highest results among the alternative configurations in the MIMIC and Memotion settings, respectively). The improvement row indicates the absolute gains of the Co-AttenDWG model over these second-best baselines, showing improvements of +0.80% in MIMIC Accuracy, +0.77% in MIMIC Macro F1, +1.69% in Memotion Accuracy, and +1.60% in Memotion Macro F1. This detailed comparison emphasizes the robustness of the Co-AttenDWG architecture and demonstrates that integrating enhanced cross-modal interactions and adaptive feature gating leads to significant performance improvements on both datasets.
@@ IV-D Ablation Study
Table III reports the results of our ablation experiments on the MIMIC and SemEval Memotion 1.0 datasets. In this study, we systematically remove specific components of the proposed Co-AttenDWG architecture to assess their impact on overall performance. The reported metrics are Test Accuracy and Macro F1, both expressed as percentages (with higher values indicating better performance). The first set of experiments involves removing critical modules one at a time. For example, the ”No ExpertFusion” experiment excludes the expert fusion module, resulting in a notable performance drop on both datasets. Similarly, when the co-attention mechanism is removed (”No Co-Attention”), the decrease in performance on both MIMIC and Memotion is evident. The ”No Cross-Attention” and ”No MambaFormer” experiments further illustrate that eliminating the cross-attention layer and the MambaFormer-based encoder modules, respectively, leads to reduced accuracy and Macro F1 scores. The removal of the convolutional operations from the MambaFormer module (”No Conv in MambaFormer”) also affects the performance, though to a slightly lesser degree. Finally, the full Co-AttenDWG model achieves the best results, demonstrating that every component contributes positively to the overall performance. These results underscore the critical role of each individual module in ensuring effective cross-modal feature integration and alignment.
@@ IV-E Findings
Table IV summarizes the performance of the proposed Co-AttenDWG architecture when various combinations of text and image backbones are employed on the MIMIC and Memotion datasets. For the MIMIC dataset, only multimodal configurations using text backbones such as mBERT and XLM-RoBERTa paired with image backbones (ResNet50, VGG16, and EfficientNetV2) are evaluated. The results reveal that the combination of XLM-RoBERTa with ResNet50 achieves the highest performance, with an Accuracy of 87.79% and an F1 score of 87.83%. This indicates that a highly expressive text encoder paired with a robust image feature extractor can effectively capture the nuances of the MIMIC dataset. For the Memotion dataset, only configurations involving BERT and DistilBERT are reported. Among these, the BERT paired with ResNet50 configuration produces the best results, yielding an Accuracy of 84.29% and an F1 score of 84.26%. These findings demonstrate that the optimal backbone combination is data-dependent, with different modalities and dataset characteristics favoring distinct model pairings. Overall, the experimental results validate the effectiveness of the Co-AttenDWG architecture in enhancing cross-modal feature integration and emphasize that careful backbone selection is crucial in achieving state-of-the-art performance in multimodal learning tasks.
@@
(c) MIMIC sample (True label: misogyny yes)
mCLIP - (✓)
VisualBERT - (✓)
Co-AttenDWG - (✓)
@@
(d) MIMIC sample (True label: misogyny yes)
mCLIP - (✓)
VisualBERT - (✗)
Co-AttenDWG - (✓)
@@
Figure 7: Case study examples on offensive content detection. The top two Figures are Memotion samples with true labels ”not_offensive” (Figure (a)) and ”hateful_offensive” (Figure (b)). In Figure (a), mCLIP fails to detect offensive content (indicated by a cross mark), while VisualBERT and Co-AttenDWG correctly classify the sample; in Figure (b), all three models correctly predict the offensive content. The bottom two Figures depict MIMIC samples with the true label ”misogyny”. In Figure (c), all models produce the correct prediction, whereas in Figure (d) VisualBERT produces an incorrect result (cross mark) while mCLIP and Co-AttenDWG classify the sample correctly.
Figure 3 shows a heatmap of the cross-attention weights in our Co-AttenDWG architecture. Each cell represents the attention strength from a text token to a particular visual feature, with warmer colors indicating higher attention and cooler colors indicating lower attention. The color scale ranges from 0 (lowest) to 1 (highest), revealing how the model dynamically aligns textual cues with relevant visual regions. Through this visualization, we can observe the degree to which certain words focus on specific parts of the image, offering insight into how the Co-AttenDWG architecture captures cross-modal dependencies and integrates complementary information. These results provide clear evidence that the learned attention distribution corresponds to meaningful semantic regions, reinforcing the effectiveness of the co-attention module. Furthermore, the heatmap demonstrates that our approach can flexibly adapt to varied input content, thereby enhancing interpretability and overall performance.
Figure 4 illustrates Grad-CAM visualizations that reveal which regions of each image the classifier deems most salient. In Figure (4(a)), the “cat meme” context is shown on the left, while the corresponding Grad-CAM heatmap on the right highlights areas where the model concentrates its attention. In Figure (4(b)), the original political meme appears on the left, and the heatmap on the right indicates the focus on both the main subject and any prominent text. These examples illustrate how the Co-AttenDWG classifier interprets visual features and aligns them with relevant textual cues to make its predictions. Furthermore, the Grad-CAM outputs provide a clear, interpretable validation of the model’s focus areas, confirming that salient image regions such as key objects and texts directly contribute to the classification outcome. This insight not only demonstrates the model’s effectiveness in capturing meaningful visual patterns but also enhances our understanding of the underlying decision-making process.
Figure 5 presents four bar plots that show the gating weights assigned to each expert for different samples in the ExpertFusion module. In Figure (5(a)), the gating mechanism assigns a relatively high weight to expert 1, indicating that for this sample the features extracted by expert 1 are the most important for the prediction, while the contributions from experts 2 and 3 are substantially lower. In Figure (5(b)), the plot displays a more even distribution of weights, with expert 2 receiving a slightly higher value than the others, which suggests that this sample benefits from a balanced combination of information from multiple experts instead of a single dominant source. In Figure (5(c)), the plot clearly shows that expert 3 receives a much higher weight compared to the other experts, indicating that for that particular sample the key discriminative features are predominantly captured by expert 3. Finally, Figure (5(d)) demonstrates another scenario in which one expert is strongly favored by the gating mechanism, with expert 1 again receiving the highest weight. Overall, these Figures provide clear insight into how the ExpertFusion module dynamically adjusts the importance of each expert in response to different input characteristics, thereby enhancing the effectiveness of the overall multi-modal representation.
IV-F Error Analysis
Figure 6 presents the confusion matrices for both the MIMIC and Memotion datasets, providing a detailed view of the classification performance. In Figure (6(a)) for the MIMIC dataset, the majority of samples are correctly classified as either Non Misogynistic or Misogynistic. However, a number of misogynistic samples are incorrectly predicted as non-misogynistic, which indicates that the model sometimes struggles with distinguishing subtle cues between these classes. These errors may stem from ambiguous visual or textual content that does not distinctly signal misogyny. In Figure (6(b)) for the Memotion dataset, we observe that misclassifications mainly occur among the offensive classes with close intensity levels, such as between ”slight” and ”very_offensive.” This suggests that there is considerable overlap in the feature space of these classes, which can be attributed to nuanced differences in language and image contexts. Overall, both Figures highlight that while the model performs well on overall classification, there remain challenges in precisely delineating fine-grained differences between closely related offensive categories. Addressing these issues may require enhanced feature extraction methods and improved annotation consistency.
Figure 7 presents four representative examples that illustrate the performance of the Co-AttenDWG architecture in detecting offensive content from memes, demonstrating its effectiveness in comparison with mCLIP and VisualBERT. In Figure (7(a)), a Memotion sample with a true label of ”not_offensive” is shown; here VisualBERT and Co-AttenDWG produce the correct outcome, while mCLIP fails to correctly classify the sample as non-offensive. Figure (7(b)) features another Memotion sample, with the true label ”hateful_offensive,” where all three models yield accurate predictions, confirming the robust performance of the models under challenging conditions. Moving to the MIMIC dataset, Figure (7(c)) illustrates a sample with a true label of ”misogyny,” and all models—mCLIP, VisualBERT, and Co-AttenDWG—successfully classify the sample, indicating consistent performance when detecting misogynistic content. In Figure (7(d)), another MIMIC sample labeled ”misogyny” is presented; while mCLIP and Co-AttenDWG correctly identify the offensive content, VisualBERT misclassifies the sample, highlighting subtle differences in model behavior. Overall, this figure provides detailed insight into the decision-making process of the examined architectures, revealing that the Co-AttenDWG model maintains high accuracy in both Memotion and MIMIC scenarios by effectively integrating multimodal cues and adapting to the unique characteristics of each input.
V Conclusions
In this work, we introduced the multi-modal Co-AttenDWG architecture, which leverages cross-modal co-attention, dimension-wise gating, dual-path encoding, and advanced expert fusion to dynamically integrate text and image features. Our extensive experiments on the MIMIC and SemEval Memotion 1.0 datasets demonstrate that our approach effectively enhances cross-modal feature integration, leading to significant improvements in classification accuracy and macro F1 scores. While our results validate the benefits of adaptive fusion in multi-modal learning, we acknowledge several limitations. First, the performance of the model is sensitive to the choice of backbone models, suggesting that optimal performance may require dataset-specific adjustments. Second, the computational overhead introduced by multiple attention and gating mechanisms remains a concern for real-time applications. Future work will focus on reducing model complexity and exploring more efficient gating strategies, as well as extending our approach to additional tasks such as retrieval and scene understanding. We also plan to investigate the integration of external knowledge sources to further improve the interpretability and robustness of the system. Overall, our findings underscore the potential of adaptive fusion techniques in multi-modal learning and open promising directions for future research.

### v2

@@ III-E Classification
IV Experiment and Result Analysis
IV-A Datasets
We evaluate our Co-AttenDWG model on two publicly available datasets that present diverse and challenging scenarios for multi-modal offensive content detection. The first dataset, SemEval-2020 Memotion Analysis 1.0 [51], is widely used as a benchmark for offensive meme classification (Offensive) on social media platforms. It captures the complex range of harmful material in meme visual and word combinations with rich annotations across four levels of offensiveness: not offensive, slightly offensive, highly offensive, and hateful offensive. The second dataset, MIMIC: Misogyny Identification in Multimodal Internet Content [52], is specifically designed to address misogynistic behavior in Hindi-English code-mixed multimodal posts, a setting that introduces unique linguistic and cultural challenges for joint text-image understanding. MIMIC comprises four distinct classification tasks targeting related yet conceptually different forms of offensive content: Misogynistic, Humiliation, Objectification, and Prejudice. Notably, the MIMIC dataset is multilingual, featuring code-mixed Hindi-English posts, which introduces additional challenges related to language diversity and code-switching in multi-modal contexts. Each of these categories requires the model to recognize subtle cues in both textual and visual modalities, making MIMIC a comprehensive testbed for evaluating fine-grained multi-label multi-modal classification in multilingual and code-mixed scenarios. The dataset is naturally imbalanced, with significant disparities in the distribution of positive and negative examples across each class. To mitigate this, we apply upsampling techniques to balance the classes, ensuring that the model receives sufficient training examples from minority classes, thereby improving generalization and robustness. Table I summarizes the original and balanced class counts for all these categories across both datasets.
Together, the Memotion and MIMIC datasets provide a rigorous evaluation framework for our architecture, enabling us to assess its capability to handle complex multi-modal inputs, code-mixed language, and subtle offensive content distinctions across both multilingual and English language content in culturally diverse contexts. This comprehensive evaluation demonstrates the effectiveness and adaptability of Co-AttenDWG in real-world multi-modal offensive content detection scenarios.
IV-B Implementation Details
We implement our Co-AttenDWG model using Python 3.12.1 and PyTorch 2.0.1 on an NVIDIA RTX 2060 GPU with 16 GB of RAM. Our optimization strategy employs the AdamW optimizer with a fixed learning rate of $2\times 10^{-5}$, training for 20 epochs while leveraging a dynamic learning rate scheduler that adjusts the rate during training for improved convergence. The model architecture is configured with 8 attention heads in the fusion modules and 4 heads in the self-attention refinement layer. The MambaFormer encoders are set with a kernel size of 3, a depth of 2 layers, and a dropout rate of 0.1 applied uniformly across all modules. All components, including the pre-trained text encoders (BERT and XLM-RoBERTa) and the image encoder (ResNet50), are fine-tuned end-to-end to maximize cross-modal feature alignment. The detailed hyperparameter settings and data preparation options are summarized in Table II.
In terms of data preparation, we carefully clean and normalize both text and images. Text is tokenized using the BERT tokenizer, while images are resized and normalized according to dataset-specific requirements. For the MIMIC dataset which contains multilingual Hindi-English code-mixed data images are resized to $\times 200$ 200 pixels to preserve detail, whereas the SemEval Memotion 1.0 dataset images are resized to $\times 160$ 160 pixels due to resource constraints. We partition both datasets into 80% training and 20% testing splits and evaluate model performance using test accuracy and macro F1-score as primary metrics. To address the inherent class imbalance particularly prominent in minority offensive categories we apply upsampling strategies to balance the class distributions, ensuring equitable representation during training. These carefully chosen hyperparameter settings and robust preprocessing techniques enable Co-AttenDWG to effectively capture fine-grained cross-modal interactions and demonstrate superior performance on complex multi-modal offensive content detection tasks.
@@ IV-B Implementation Details
(a) Sample 1: Balanced gate, slightly favors image; moderate confidence.
(b) Sample 2: Low confidence, gate favors text.
(c) Sample 3: High confidence, strong image gating.
(d) Sample 4: Confident prediction, attention sharp on image.
Figure 6: Self-attention heatmap visualizations for selected samples from the fine-grained interpretability analysis. Each subfigure illustrates how the model distributes attention between modalities and across sequence/image regions for the given example.
@@ IV-C Baseline Comparison
Table III presents an extensive performance comparison between our proposed Co-AttenDWG model and a wide range of baseline and state-of-the-art multimodal models evaluated on the MIMIC and SemEval Memotion datasets. The evaluation includes key offensive content detection categories such as Misogyny, Objectification, Prejudice, Humiliation, and Offensive Content from the Memotion dataset, reporting both accuracy and F1 scores to comprehensively capture performance. Co-AttenDWG consistently outperforms all baseline models across the majority of categories, establishing new state-of-the-art results with accuracy and F1 scores of 87.19% and 87.16% for Misogyny detection, 94.80% for both metrics in Objectification, 97.15% in Prejudice, and 84.29% accuracy alongside 84.26% F1 on the Memotion offensive content detection task. These results demonstrate Co-AttenDWG’s exceptional ability to capture subtle and complex cross-modal interactions between textual and visual modalities, crucial for nuanced offensive content understanding. In the Humiliation category, VisualBERT emerges as the highest-performing baseline, achieving 98.91% in both accuracy and F1 scores. Co-AttenDWG closely follows with a very competitive 98.80%, trailing by a marginal 0.11 percentage points. Notably, other strong vision-language models such as mCLIP, ALBEF, and BLIP also perform well in this category, with mCLIP and VisualBERT setting a high bar for multi-modal understanding. Despite the slight dip in the Humiliation metric, Co-AttenDWG surpasses these models substantially in all other categories, reflecting its balanced and robust performance profile. The inference time of Co-AttenDWG is measured at 31.1 milliseconds per sample on an NVIDIA RTX 2060 12GB GPU with a batch size of one. This is competitive considering the advanced architectural components, such as dual-path encoders, co-attention with dimension-wise gating, and expert fusion modules, which collectively contribute to its superior accuracy and fine-grained modeling capacity.
Overall, these results underscore the strength of Co-AttenDWG in effectively integrating and refining multi-modal signals for offensive content detection, outperforming current vision-language models and baseline architectures. The model’s ability to generalize across diverse and culturally nuanced datasets like MIMIC and Memotion highlights its potential applicability in real-world scenarios requiring sensitive and precise offensive content moderation.
@@ IV-D Ablation Study
Table IV presents a comprehensive combinatorial ablation study evaluating the impact of removing or modifying key components of the Co-AttenDWG model on performance across the MMIC and Memotion datasets. Each variant disables or alters one or more major modules, such as ExpertFusion (EF), Co-Attention (CA), Cross-Attention (XA), MambaFormer (MF), and Fine-grained Fusion (FF), or reduces the number of attention heads. The results show consistent performance degradation across all tasks when any of these components are removed, confirming their individual contributions to the overall model effectiveness. Notably, the full Co-AttenDWG model achieves the highest accuracies and F1 scores, reaching up to 87.19% accuracy on Misogyny and 98.80% on Humiliation, demonstrating robust multi-modal learning capabilities. The largest performance drops occur when multiple critical modules are removed simultaneously, such as the combination of ExpertFusion, Co-Attention, and MambaFormer, which substantially lowers results across all evaluated categories. This ablation analysis validates the importance of each architectural element in capturing fine-grained, cross-modal interactions essential for state-of-the-art offensive content detection and sentiment understanding.
@@ IV-E Findings
Table V presents an ablation study analyzing the impact of core architectural hyperparameters on the Co-AttenDWG model’s performance across multiple labels in the MIMIC and Memotion validation sets. The hyperparameters explored include the number of experts, cross-attention heads, co-attention heads, MambaFormer kernel size and depth, dropout rate, input image resolution (pixel value), and learning rate. The results demonstrate that increasing the number of experts and attention heads generally enhances performance, with the best configuration employing 8 experts, 8 cross-attention heads, and 4 co-attention heads. A smaller kernel size of 3 and a shallow MambaFormer depth of 2 layers, combined with a moderate dropout rate of 0.10, also contribute to optimal results. Image resolution plays a role, with the best setting utilizing a combination of 200 and 160 pixels depending on the dataset. Learning rate tuning around $2\times 10^{-5}$ further stabilizes training. This optimal setting yields the highest macro F1 scores across all labels: 87.16% for Misogyny, 94.80% for Objectification, 97.15% for Prejudice, 98.80% for Humiliation, and 84.26% for the Memotion dataset. These findings underscore the importance of carefully balancing architectural complexity and regularization to maximize multi-label multimodal classification performance.
Table VI presents an ablation study evaluating the impact of different backbone combinations on the performance of the Co-AttenDWG model across the MIMIC and Memotion datasets. We explore three text encoders (mBERT, XLM-RoBERTa, and BERT variants) paired with three image backbones (ResNet50, VGG16, and EfficientNetV2) to assess how backbone selection affects multi-modal offensive content detection. Across all MIMIC sub-tasks, including Misogyny, Objectification, Prejudice, and Humiliation, the results show that the multilingual XLM-RoBERTa text encoder together with the ResNet50 image backbone achieves the best overall performance, with accuracies and F1 scores consistently above 87% and 94%, respectively. Notably, this combination also maintains high effectiveness on the Memotion dataset. While mBERT backbones also deliver strong results, particularly with ResNet50 and VGG16, they generally perform slightly below the top-performing XLM-RoBERTa models. EfficientNetV2 backbones show comparatively lower results, especially in the Humiliation category. The BERT and DistilBERT variants, though effective on Memotion, lack reported results for several MIMIC sub-tasks, reflecting possible limitations in handling code-mixed multilingual data. These findings underline the critical role of backbone selection in multi-modal architectures and support the use of powerful, multilingual text encoders and strong visual backbones to maximize performance in complex offensive content detection tasks.
TABLE VII: True Negatives (TN), False Positives (FP), False Negatives (FN), and True Positives (TP) for all binary and multiclass tasks.
Task / Class TN FP FN TP
Humiliation 923 10 0 882
Misogyny 455 49 85 410
Objectification 639 41 34 671
Prejudice 763 64 16 770
Memotion (Offensive Content)
not_offensive 1446 150 148 382
slight 1462 114 172 378
very_offensive 1492 121 80 433
hateful_offensive 1578 15 0 533
@@
(c) MIMIC sample with true labels: misogyny, prejudice, objectification, humiliation.
mCLIP - misogyny (✓), prejudice (✓), objectification (✗), humiliation (✓).
VisualBERT - misogyny (✓), prejudice (✓), objectification (✓), humiliation (✓).
ALBEF - misogyny (✓), prejudice (✗), objectification (✓), humiliation (✓).
BLIP - misogyny (✓), prejudice (✗), objectification (✓), humiliation (✓).
Co-AttenDWG - misogyny (✓), prejudice (✓), objectification (✓), humiliation (✓).
@@
(d) MIMIC sample with true labels: misogyny, prejudice, objectification, humiliation.
mCLIP - misogyny (✓), prejudice (✗), objectification (✗), humiliation (✓).
VisualBERT - misogyny (✗), prejudice (✗), objectification (✓), humiliation (✓).
ALBEF - misogyny (✗), prejudice (✓), objectification (✓), humiliation (✓).
BLIP - misogyny (✓), prejudice (✓), objectification (✓), humiliation (✓).
Co-AttenDWG - misogyny (✓), prejudice (✓), objectification (✓), humiliation (✓).
@@
Figure 7: Case study examples on offensive content detection. The top two subfigures show Memotion samples with true labels ”not_offensive” (a) and ”hateful_offensive” (b). In (a), mCLIP fails to detect offensive content (✗), while VisualBERT, ALBEF, BLIP, and Co-AttenDWG correctly classify the sample (✓); in (b), all models correctly predict the offensive content (✓). The other two subfigures depict MIMIC samples with true labels including misogyny, prejudice, objectification, and humiliation. In (c), all models correctly classify all labels (✓), whereas in (d), VisualBERT incorrectly classifies misogyny and prejudice (✗) but correctly identifies objectification and humiliation (✓), while mCLIP, ALBEF, BLIP, and Co-AttenDWG correctly classify all labels (✓).
IV-F Interpretability
Figure 3 presents a detailed heatmap visualization of the cross-attention weights learned by the Co-AttenDWG architecture. Each cell in the heatmap quantifies the attention strength from a specific textual token to a corresponding visual feature, with warmer colors signifying higher attention values and cooler colors indicating lower values on a normalized scale from 0 to 1. This visualization demonstrates how the model dynamically aligns relevant textual cues with semantically meaningful regions within the image, effectively capturing intricate cross-modal dependencies. Complementing this, Figure 4 shows Grad-CAM visualizations that highlight salient image regions influencing the classifier’s predictions. Specifically, subfigure (4(a)) depicts a “cat meme” where the heatmap emphasizes key visual elements aligned with textual content, whereas subfigure (4(b)) illustrates a political meme, indicating attention over both the principal subject and embedded textual information. Collectively, these visualizations validate the model’s capacity to interpret and fuse multi-modal features in a meaningful and interpretable manner. Furthermore, Figure 5 presents bar plots that illustrate the gating weights assigned by the ExpertFusion module to different experts across diverse samples. The shift in gating distributions from a preponderance of one expert to more balanced weightings among several experts demonstrates the module’s adaptability and flexibility in adjusting the impact of different feature extractors according to input data. Such dynamic expert weighting is necessary to improve the model’s overall representational capability and integrate complementary multi-modal information in an efficient manner.
Figure 6 presents an in-depth examination of the fine-grained interpretability of the Co-AttenDWG model through a series of self-attention heatmap visualizations for four representative test samples. Subfigure 6(a) (Sample 1) shows the gating network allocating balanced attention weights to both text and image modalities, with a slight emphasis on visual features, reflecting moderate confidence and demonstrating the model’s capacity to integrate complementary signals harmoniously. Subfigure 6(b) (Sample 2) illustrates a low-confidence instance where the gating mechanism predominantly favors the textual modality, indicating that the model appropriately relies more heavily on language features when visual cues are ambiguous or less informative. Subfigure 6(c) (Sample 3) depicts a high-confidence prediction characterized by a strong gating bias toward the image modality, highlighting the model’s ability to prioritize salient visual information when it serves as a more definitive classification indicator. Lastly, Subfigure 6(d) (Sample 4) demonstrates a confident prediction accompanied by sharply focused self-attention on specific spatial regions within the image, evidencing the model’s proficiency in localizing and attending to critical visual cues that substantively contribute to its decision-making process. Together, these subfigures reveal the dynamic, context-sensitive fusion strategy employed by Co-AttenDWG, illustrating how the interplay between gating and attention mechanisms adapts to the input data to improve interpretability and classification robustness in complex multi-modal scenarios.
IV-G Error Analysis
Table VII presents detailed counts of true negatives (TN), false positives (FP), false negatives (FN), and true positives (TP) for both the MIMIC and Memotion datasets, providing an in-depth analysis of classification performance. For the MIMIC dataset, the Humiliation class exhibits excellent classification performance, with 923 true negatives and 882 true positives, and minimal errors, indicating the model’s strong ability to identify this category. In contrast, the Misogyny class shows higher confusion, with 49 false positives and 85 false negatives, highlighting challenges in correctly distinguishing misogynistic content due to subtle textual and visual cues. Objectification and Prejudice classes demonstrate moderate misclassification rates, with a balanced distribution of false positives and false negatives, suggesting overlapping features in multimodal inputs contribute to classification ambiguity. For the Memotion dataset, which focuses on offensive content intensity levels, the “not_offensive” and “hateful_offensive” categories show relatively strong separability with higher true negatives and true positives and fewer misclassifications. However, intermediate classes such as “slight” and “very_offensive” have notable misclassification rates, reflecting the difficulty in distinguishing nuanced differences between similar offensive intensities. These observations suggest that while the model effectively handles broad category distinctions, fine-grained discrimination among closely related classes remains challenging, underscoring the potential for improved feature extraction and more consistent annotation to enhance multi-class classification performance.
Figure 7 presents four illustrative examples demonstrating the efficacy of the Co-AttenDWG architecture for offensive content detection in multimodal memes, alongside comparisons with mCLIP and VisualBERT. In Figure (7(a)), a Memotion sample labeled as ”not_offensive” is examined, where both VisualBERT and Co-AttenDWG correctly classify the sample, whereas mCLIP fails to detect its non-offensive nature. This indicates that Co-AttenDWG and VisualBERT possess greater sensitivity to subtle non-offensive cues in multimodal content. In contrast, Figure (7(b)) shows a ”hateful_offensive” Memotion meme that all three models classify correctly, reflecting their robustness when identifying clearly offensive content. Shifting focus to the MIMIC dataset, Figure (7(c)) depicts a misogynistic post that all models accurately recognize, signifying consistent detection capabilities for misogyny across architectures. However, in Figure (7(d)), another misogynistic MIMIC example reveals divergent model behaviors: while mCLIP and Co-AttenDWG correctly identify the offensive content, VisualBERT misclassifies the instance, underscoring Co-AttenDWG’s improved generalization in challenging or ambiguous cases. Collectively, these examples elucidate how Co-AttenDWG effectively integrates textual and visual information to maintain high classification accuracy across diverse, real-world scenarios, surpassing or matching state-of-the-art alternatives in both non-offensive and offensive content detection.
V Limitations and Future Works
While the Co-AttenDWG architecture achieves notable improvements in multimodal offensive content detection, several limitations remain. First, although the model effectively fuses textual and visual modalities, it can face challenges when processing highly ambiguous, context-dependent content, especially within code-mixed or low-resource language scenarios. The use of fixed pre-trained backbones such as BERT and ResNet may limit adaptability to emerging linguistic patterns and novel visual meme formats, potentially impacting robustness over time. Second, the current model primarily targets single-label or multi-class classification tasks and does not explicitly model hierarchical or multi-label dependencies where overlapping offensive content categories coexist. This restricts the model’s ability to fully capture the nuanced relationships among different forms of offense. Third, while upsampling addresses class imbalance, it may inadvertently cause overfitting or bias toward synthetic samples, underscoring the need for more sophisticated imbalance mitigation strategies, including focal loss or data augmentation approaches.
For future work, incorporating Vision Transformer (ViT) models could significantly enhance visual representation learning due to their superior ability to capture long-range dependencies and global contextual information in images. Extending the architecture to handle additional modalities such as audio and video would further broaden applicability in multimedia-rich social platforms. Furthermore, developing dynamic fusion techniques that adaptively modulate interactions based on input complexity, as well as incorporating continual learning paradigms, can improve model generalization over time and across domains. The model could also be adapted and evaluated on other multimodal tasks such as sentiment analysis, hate speech detection, or misinformation classification, to validate its generalizability and robustness across diverse applications. Finally, integrating explainability mechanisms will be crucial to increase transparency, foster trust, and support ethical deployment in real-world content moderation systems.
VI Conclusions
In this work, we proposed Co-AttenDWG, a novel multimodal architecture that effectively integrates textual and visual information through dual-path encoding, co-attention with dimension-wise gating, and expert fusion. Our approach dynamically captures fine-grained cross-modal interactions, enabling robust alignment of heterogeneous features. Extensive experiments on the MIMIC and SemEval Memotion datasets demonstrated that Co-AttenDWG consistently outperforms state-of-the-art baselines, achieving superior accuracy and F1 scores across multiple offensive content detection categories. The qualitative analyses further reveal the model’s ability to focus on semantically meaningful regions in both modalities, highlighting its interpretability and adaptability to diverse, culturally nuanced contexts. While challenges remain in handling ambiguous content and class imbalance, our results establish a strong foundation for future multimodal research. The proposed framework can be extended to incorporate advanced visual encoders such as Vision Transformers and to support richer multimodal inputs beyond text and images. Overall, Co-AttenDWG advances the field of multimodal offensive content detection by providing a powerful, flexible, and interpretable solution capable of addressing complex real-world scenarios.